OpenStudy (tatumlee):

(6x^-2)^2(0.5x)^4 how can i solve this?

OpenStudy (thesmartestderpalive):

Do you want a calculator that can solve this? Or do you want to do it manually?

OpenStudy (tatumlee):

Which is better?

OpenStudy (thesmartestderpalive):

I'd say the calculator

OpenStudy (thesmartestderpalive):

Here: www.pathway.com

OpenStudy (thesmartestderpalive):

Mathway*

OpenStudy (tatumlee):

I have to show my work on this, will the calculator interfere with that?

OpenStudy (zale101):

Do you know all the rules for exponents?

OpenStudy (tatumlee):

I know some of them I am not sure if I know them all

OpenStudy (zale101):

This is the rule that you need to apply to your question \(\large (x^m)^n)=x^{m*n}\)

OpenStudy (zale101):

For example, if you had \large \((2x^2)^3)=8x^{2*3}=8x^{6}\)

OpenStudy (tatumlee):

Is there a name for that rule?

OpenStudy (zale101):

It's he Power Rule for exponents.

OpenStudy (zale101):

the*

OpenStudy (tatumlee):

@TheSmartestDerpAlive

OpenStudy (tatumlee):

I am confused, aren't there more steps than that?